Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4c)
tomelleri - 23:01 Monday 15 September 2025 (67709) Print this report
Operator Report - Afternoon shift

ITF found locked in LOW_NOISE_3 with SCIENCE mode; BNS Range ~55 Mpc. ITF set to CALIBRATION mode at 15:30UTC for routine calibration activities (Calibrations with ITF locked, ENV injections, ISC injections). ITF in COMMISSIONING mode at 16:13UTC required for injections via ENV_MAIN node. ITF unlocked due to earthquake (id us7000qwch, see pdf 1) at 17:26UTC. Automation relocked to LN3 on first attempt at 18:09UTC, back to SCIENCE mode with Autoscience engaged at 18:36UTC in order to finish LSC injection. ITF left locked

Guard tour (UTC)
17:58 -> 18:33

Sub-system reports

Calibration
(times in UTC)
15:30 - 15:49 CALIBRATED_DF_SENSITIVITY
15:50 - 16:09 CALIBRATED_DF_PCAL
18:16 - 18:33 run inject_lsc.py

EnvMoni
16:13 - 17:08UTC STARTING_NOISE_INJECTIONS. Both magnetic and acoustic injections performed thanks to Tringali's help, close supervision and loudspeaker PDU bypass (entry #67562)
Acoustic injections performed under the effects of earthquake (id us7000qwch) visible as excess noise on all F0_Z loops from DMS.

Images attached to this report
Non-image files attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×